# Identifying noise transients in gravitational-wave data arising from nonlinear couplings

### v1

@@ 2.1 Usefulness of the bicoherence: An example
Figure 2: The two circles in the top-left bicoherence map mark the two phase-coupled (PC) bifrequencies, namely, (20,80) Hz and (20,100) Hz, corresponding to the signal in Eq. (2.1), with bicoherence of $\simeq 1$. Contrastingly, in the bottom-left bicoherence map of the corresponding NPC signal, the same region shows a bicoherence of $\simeq 0$, implying lack of phase coupling. Phase randomization (PR), which was used here, improves the identification of PC bifrequencies. To analyze the behavior of bicoherence at and around the bifrequencies $(20,80)$ Hz and $(20,100)$ Hz, line plots are shown in the middle two panels, with $f_{l}\in[70-110]$ Hz and $f_{k}$ kept fixed at 20 Hz. The vertical dashed black lines correspond to the bifrequencies (20,80) Hz and (20,100) Hz. Top-middle: At the phase-coupled bifrequencies, the bicoherence is close to one. The bicoherence at those bifrequencies dips to low values for the non-phase-coupled case (bottom-middle plot). Note that for this (NPC) case, we chose $\Phi_{\rm zp}=0=\Phi_{\rm zn}$. The right panel shows the bicoherence histogram of the left panel. The bottom-right figure shows that the probability of misclassifying an NPC as PC signal is 11.76% for $b_{\rm pr}>0.995$, but drops to 0% above a value of $b_{\rm pr}=0.996$ (in this experiment).
@@
To test the above expectations through simulations, we added the PC and NPC versions of $z(t)$, separately, into 2000 realizations of noise, $n(t)$. For these simulated data sets, for PC and NPC cases, we then computed the (auto-)bicoherence values and their corresponding biphase variances without applying phase randomization. We repeated the experiment after applying phase randomization, and plotted results from both studies in Fig. 1, which shows that phase randomization improves the distinguishability of phase-coupled signals from non-phase-coupled signals.
@@
The bicoherence maps of the same simulated signals, for the PC and NPC cases, are shown in Fig. 2 (left panel). The right panel in Fig. 2 is the non-redundant  [^1] See A. bifrequencies histogram. The middle panel shows the values of $b_{\rm pr}$ as a function of $f_{l}$ when $f_{k}$ is held fixed at 20 Hz. The bottom-right figure shows that the probability of misclassifying an NPC as PC signal is 11.76% when the bicoherence ($b_{\rm pr}$) threshold is set to 0.995, but drops to 0% for $b_{\rm pr}>0.996$ (in this experiment). The non-linear phase couplings in the signal are identified correctly through the phase randomized bicoherence method. In this paper, the bicoherence maps are positive quadrants of the bicoherence plane (see A). Note that each octant in the bicoherence plane captures complete information about the bifrequencies owing to its eight-fold symmetry.
@@ 3 Estimating detection threshold
In the previous section, we demonstrated that applying phase randomization in the computation of bicoherence improves the distinguishability of PC and NPC cases. This improvement rests on the following reasonable assumptions about the signals: (i) the phase component of bifrequencies $\Phi_{k}^{\alpha}$ and $\Phi_{l}^{\alpha}$ are uniformly distributed random variables, over $(-\pi,\pi]$; (ii.a) For the NPC bifrequencies, we shall take that $\Phi_{m}^{\alpha}\neq 0\neq\Phi_{\rm d}^{\alpha}$; (ii.b) For NPC cases, $\Phi_{m}^{\alpha}$ and $\Phi_{\rm d}^{\alpha}$ are uniformly distributed as $\sim U(-\pi,\pi]$ [25] and independent of $\Phi_{k}^{\alpha}$ and $\Phi_{l}^{\alpha}$; (iii) For the PC bifrequencies, $\Phi_{\rm d}=0$ – although, in practice, it is not exactly zero and can be assumed to have a normal distribution, i.e., $\Phi_{\rm d}\sim N(0,\sigma_{\Phi}^{2})$ [27]. However, the true phase-coupled bifrequencies can become indistinguishable from noise when the signal’s power or the strength of the phase coupling is low. This can lead to a false dismissal of PC. Therefore, a test is required to decide whether or not there is phase coupling present at any bifrequency. We assume two hypotheses that are tested on the simulated data:
@@ 4 Application to gravitational-wave data
Importantly, failing to apply artificial phase randomization in the computation of auto-bicoherence can lead to false detection of PC bifrequencies. Consequently, the bicoherence maps also show different patterns depending on whether PR is used or not. One of the blip glitches that occurred in a LIGO detector during the O2 observation run with high excess power is demonstrated in Figure 4, where the effect of PR can be clearly discerned. We continue to apply PR in all our bicoherence analysis below.
Figure 4: The bicoherence map of a blip glitch, constructed without phase randomization (PR), is shown on the left. The colorbar shows the bicoherence. It incorrectly shows many coupled bifrequencies, which disappear when the bicoherence is computed using phase randomization, as shown in the middle figure. This figure also reveals that the glitch may have been triggered from disturbances in frequencies that up-converted due to phase couplings. Column-3: For comparison, we show here the PR bicoherence map of the background noise data that does not contain any glitch. Such data show the absence of any significant phase coupling.
@@ 4.1 Using bicoherence to classify glitches
Instances of transient excess noise often show up in auxiliary channels and are detected by tools such as Omicron [31]. This tool performs a multi-resolution time-frequency analysis of detector data around such instances. Glitches with excess power above a given threshold are flagged for further inspection in data quality studies and astrophysical searches. Also, the time-frequency (TF) maps or spectrograms of Omicron serve to classify glitches via human intervention and machine learning  [29, 32].
@@
Figure 5: A comparison of second-order and third-order statistics of glitches. The top row shows the spectrograms of pairs of blips (columns 1-2), koi fish (columns 3-4) and “extremely loud” glitches (columns 5-6). The spectrograms look similar within each category, e.g., blips, simply because the glitches were categorized by the patterns revealed in those maps in the first place. They are different – to varying degrees – across categories, e.g., blips and extremely loud glitches. The bottom row shows the corresponding phase-randomized bicoherence maps, constructed using 4 sec data chunks centered at the glitches. Variations in the bicoherence maps within some glitch categories (e.g., columns 3 and 4 of koi fish) suggest the presence of different phase-coupled bifrequencies or a difference in the nature of that coupling. On the other hand, similarities in these maps across glitch categories (e.g., columns 1 and 4) suggest the presence of similar nonlinear phase couplings in different glitch types. Information in both types of maps, combined, can help in troubleshooting the origin of these noise artifacts.
@@
The bicoherence maps of the blip, koi fish, and extremely loud glitches are produced by applying the phase-randomized auto-bicoherence analysis to look for $3^{\rm rd}$-order similarities in each class. If the observed glitches have a common source, their bicoherence maps will have similarities owing to common bifrequencies related to that source. Such expectations have been borne out in the study of $2^{\rm nd}$-order statistics, such as the TF maps of Omicron, as depicted in the upper panel of Fig. 5. The top row in that figure shows the spectrograms of pairs of blips (columns 1-2), koi fish (columns 3-4) and “extremely loud” glitches (columns 5-6). The spectrograms look similar within each category, e.g., blips, simply because the glitches were categorized by the patterns revealed in those maps. They are different – to varying degrees – across categories (e.g., blips and extremely loud glitches). The bottom row shows the corresponding phase-randomized bicoherence maps, constructed using 4sec data chunks centered at the glitches. Variations in the bicoherence maps within some glitch categories (e.g., columns 3 and 4 of koi fish) suggest the presence of different phase-coupled bifrequencies or a difference in the nature of that coupling. On the other hand, similarities in these maps across glitch categories (e.g., columns 1 and 4) suggest the presence of similar nonlinear phase couplings in different glitch types. These maps indicate, e.g., that all blips are not made identically. In future, cross-bicoherence maps can help diagnose their origins by revealing what detector systems couple non-linearly to produce them.
It is also interesting to inquire if bicoherence maps can be used to distinguish between Compact Binary Coalescence (CBC) signals and glitches arising from non-linear couplings. We leave this question for a future work to investigate but provide some preliminary findings here. By applying the PR auto--bicoherence analysis to GWTC-3 [7] catalog events we examine the apparent phase coherence in the data with CBC signals. We noticed significant phase coherence in a few GW events. Two of those, namely, GW events GW150914 [36], and GW190521_074359 [5] are presented in Fig. 6 along with a simulated signal for aiding the interpretation of their bicoherence maps. The presence of significant bicoherence is evident there. To prepare the GW strain data for the bicoherence analysis, first a high-pass filter is applied with cut-off frequency 20 Hz, followed by whitening and the application of a bandpass filter, from 30 to 350 Hz, with notches of the first three power line harmonics [37]. The noise power-spectral density (PSD) is estimated from 32 sec of data using the Welch method.
To ensure that the phase couplings revealed in GW events’ bicoherence maps are not from glitches, we modeled two signals similar to the GW150914 event, with and without strain noise. In Fig. 6 (row 1, column 3), the 8 second long time-series noise was simulated by using the H1 noise PSD available in Ref. [38] for the GW190514 event. A simulated CBC signal was added to that noise before producing the plot. On the other hand, in Fig. 6 (row 2, column 3) no noise was added to the CBC signal. The bicoherence patterns in the map of the modeled signal, with added simulated noise, resembles those of the GW150914 signal, as shown in Fig. 6 (row 1, column 3). These signals were produced from the SEOBNRv4_opt waveform [39], with component masses of 36 Mdirect-$\odot$ and 30 Mdirect-$\odot$. This family of waveforms was constructed in the literature by calibrating effective one-body waveforms with an extensive set of numerical relativity waveforms, and have been employed in LIGO-Virgo CBC searches (see, e.g., Ref. [40]. The matched-filter SNR of the signal is 20. At the very least, the knowledge of these bicoherence patterns will be important in not confusing real CBC signals for noise artifacts and, therefore, for the purpose of veto safety.
@@ 4.2 Recurrent bifrequencies in the quieter Omicron tiggers
We also report that the nonlinear couplings between the mechanical resonances of test masses reported in our earlier work [18] are found to exist in more recent observation runs as well. Specifically, the bounce and the violin modes of test masses (which register in the bicoherence map at bifrequency (10, 495.2) Hz) as well as the power line harmonics and their sub-harmonics identified in that analysis can be found in O3 too.
@@
198.4, 303.2
501.6
501.61
Violin mode 1st harmonic ITMY Mode 1
501.68
Violin mode 1st harmonic ITMY Mode 2
@@
10.0, 383.2
393.2
393.2
Calibration line non-linearity
16.4, 477.2
493.6
493.61514
Likely calibration line mixing
@@
30.0, 54.4
84.4
84.4
Calibration line non-linearity
38.0, 66.4
104.4
104.4
40.0, 448.8
488.8
488.78625
Likely calibration line mixing
48.8, 250.8
299.6
299.6
Beam splitter violin mode region
@@
68.4, 420.4
488.8
488.78625
Likely calibration line mixing
78.0, 423.6
501.6
501.63223
Violin mode 1st harmonic
@@
134.0, 337.6
471.6
471.68625
Likely calibration line mixing
150.8, 242.4
393.2
393.2
Calibration line non-linearity
155.6, 333.2
488.6
488.78625
Likely calibration line mixing
206.0, 282.4
488.4
488.48624
221.2, 287.6
508.8
508.84515
Violin mode 1st harmonic
Table 2: A list of bifrequencies with $b_{\rm pr}\geq 0.98$. The sum of the two frequencies in a significant bifrequency pair is compared with the frequency of the gated lines in the O2 and O3 observation runs. The closest match of those frequencies is reported here (in H1). No such match was found in L1 at the same level of significance.

### v2

@@ 2.1 Usefulness of the bicoherence: An example
Figure 2: The two circles in the top-left bicoherence map mark the two phase-coupled (PC) bifrequencies, namely, (20,80) Hz and (20,100) Hz, corresponding to the signal in Eq. (2.1), with bicoherence of $\simeq 1$. Contrastingly, in the bottom-left bicoherence map of the corresponding NPC signal, the same region shows a bicoherence of $\simeq 0$, implying lack of phase coupling. Phase randomization (PR), which was used here, improves the identification of PC bifrequencies. To analyze the behavior of bicoherence at and around the bifrequencies $(20,80)$ Hz and $(20,100)$ Hz, line plots are shown in the middle two panels, with $f_{l}\in[70-110]$ Hz and $f_{k}$ kept fixed at 20 Hz. The vertical dashed black lines correspond to the bifrequencies (20,80) Hz and (20,100) Hz. Top-middle: At the phase-coupled bifrequencies, the bicoherence is close to one. The bicoherence at those bifrequencies dips to low values for the non-phase-coupled case (bottom-middle plot). Note that for this (NPC) case, we chose $\Phi_{\rm zp}=0=\Phi_{\rm zn}$. The right panel shows the bicoherence histogram of the left panel. The bottom-right figure shows that the probability of misclassifying an NPC as PC signal is 7e-4 for $b_{\rm pr}>0.995$, but drops to 0% above a value of $b_{\rm pr}=0.996$ (in this experiment).
@@
Given the importance of the bicoherence maps in this work, it is worthwhile to go over this logic in more detail: In the example of Eq. (2.1), $z(t)$, or its Fourier transform $\tilde{Z}$, has power in four frequencies, namely, $802f_{x}=$ 80 Hz, $1202f_{y}=$ 120 Hz, $f_{\rm zp}=(60+40){\;\rm Hz}=100$ Hz and $-f_{\rm zn}=(60-40){\;\rm Hz}=20$ Hz (keeping to positive frequency values). This will remain true for all $M$ segments constructed from the original time-series $z(t)$ by following the procedure dsrcribed above. As can be established by setting $\tilde{X}_{\alpha}=\tilde{Y}_{\alpha}\equiv\tilde{Z}_{\alpha}$, for all $\alpha$, in the definition of the bispectrum in Eq. (5), one of the three conditions that is necessary but not sufficient for it, and the associated auto-bicoherence, to assume large values for any bifrequency is that $f_{m}=f_{l}+f_{k}$. For the current example, $f_{l}$ and $f_{k}$ can be any pair chosen from the same set of four frequencies, viz., $\{20,80,100,120\}$ Hz, but the bispectrum will be large only when $f_{m}$ also assumes one of the same four values. This is the second necessary condition. Consequently, when $f_{l}=20{\;\rm Hz}=f_{k}$, one has $f_{m}=40$ Hz, which, however, is not present in the above set of four frequencies and, hence, one does not find strong bispectrum, or bicoherence, at the bifrequency of (20,20) Hz. For similar reasons, the bifrequency of (80,80) Hz also does not have a strong bicoherence value.
Contrastingly, with $f_{l}=20$ Hz and $f_{k}=80$ Hz, one has $f_{m}=100$ Hz, which is present in the set of four. Hence, the bifrequency (20,80) Hz satisfies the first two conditions. The third and final condition that must be satisfied is that $\Phi_{\rm d}^{\alpha}\equiv\Phi^{\alpha}(f_{k})+\Phi^{\alpha}(f_{l})-\Phi^{\alpha}(f_{m})$, defined above Eq. (4), must vanish at those frequencies, for each one of the $M$ segments. In this example, the phase-coupled signal obeys, for all segments, $\Phi_{\rm d}^{\alpha}=\Phi^{\alpha}({\rm 20~{}Hz})+\Phi^{\alpha}({\rm 80~{}Hz})-\Phi^{\alpha}({\rm 100~{}Hz})=-\Phi_{\rm zn}+2\Phi_{x}-\Phi_{\rm zp}=-(\Phi_{x}-\Phi_{y})+2\Phi_{x}-(\Phi_{x}+\Phi_{y})$, which vanishes regardless of the specific values that $\Phi_{x}$ and $\Phi_{y}$ may take in a given realization of the (quadratic) phase-coupling in Eq. (2.1). Thus, the signal $z(t)$ in this example obeys all three conditions for the bifrequency (20,80) Hz and, consequently, its bicoherence assumes a near unity value at that bifrequency. Note that the noise $n(t)$ will also contribute a random element to $\Phi_{\rm d}^{\alpha}$. However, when taking the average over a sufficiently large number of segments $M$ its influence will be small. For true phase-coupled signals present, like in Eq. (2.1), their contribution to the biphase $\Phi_{\rm d}^{\alpha}$ will be consistently zero for all $M$ segments, which helps the bicoherence attain a value close to unity.
Similarly, with $f_{l}=20$ Hz and $f_{k}=100$ Hz, one has $f_{m}=120$ Hz, which is present in the set of four. Hence, it too receives strong bispectrum and bicoherence values. Owing to the symmetry properties of the bispectrum discussed earlier the bifrequencies (80,20) Hz and (100,20) Hz will also receive large values for those statistics. Note, however, that bispectrum / bicoherence maps often do not display these regions of redundant bifrequencies (see A for more details). One can likewise show that the remaining bifrequency pairs constructed from the aforementioned set of four frequencies will not receive significant bicoherence values.
@@
To test the above expectations about the bicoherence results for the above example through simulations, we added the PC and NPC versions of $z(t)$, separately, into 2000 realizations of noise, $n(t)$. For these simulated data sets, for PC and NPC cases, we then computed the (auto-)bicoherence values and their corresponding biphase variances without applying phase randomization. We repeated the experiment after applying phase randomization, and plotted results from both studies in Fig. 1, which shows that phase randomization improves the distinguishability of phase-coupled signals from non-phase-coupled signals.
@@
The bicoherence maps of the same simulated signals, for the PC and NPC cases, are shown in Fig. 2 (left panel). The right panel in Fig. 2 is the non-redundant  [^1] See A. bifrequencies histogram. The middle panel shows the values of $b_{\rm pr}$ as a function of $f_{l}$ when $f_{k}$ is held fixed at 20 Hz. The bottom-right figure shows that the probability of misclassifying an NPC as PC signal is 7e-4 when the bicoherence ($b_{\rm pr}$) threshold is set to 0.995, but drops to 0% for $b_{\rm pr}>0.996$ (in this experiment). The non-linear phase couplings in the signal are identified correctly through the phase randomized bicoherence method. In this paper, the bicoherence maps are positive quadrants of the bicoherence plane (see A). Note that each octant in the bicoherence plane captures complete information about the bifrequencies owing to its eight-fold symmetry.
@@ 3 Estimating detection threshold
In the previous section, we demonstrated that applying phase randomization in the computation of bicoherence improves the distinguishability of PC and NPC cases. This improvement rests on the following reasonable assumptions about the signals: (i) the phase component of bifrequencies, $\Phi_{k}^{\alpha}$ and $\Phi_{l}^{\alpha}$, are uniformly distributed random variables, over $(-\pi,\pi]$; (ii.a) For the NPC bifrequencies, we shall take that $\Phi_{m}^{\alpha}\neq 0\neq\Phi_{\rm d}^{\alpha}$; (ii.b) For NPC cases, $\Phi_{m}^{\alpha}$ and $\Phi_{\rm d}^{\alpha}$ are uniformly distributed as $\sim U(-\pi,\pi]$ [30] and independent of $\Phi_{k}^{\alpha}$ and $\Phi_{l}^{\alpha}$. The worst case scenario, as discussed above, is when all $\Phi_{d}^{\alpha}$ are identical, say, $\Phi_{d}$, which falsely produces large bicoherence values and necessitated the use of phase randomization; (iii) For the PC bifrequencies, $\Phi_{\rm d}^{\alpha}=0$ for all $\alpha$ – although, in practice, they are not exactly zero and can be assumed to have a normal distribution, i.e., $\Phi_{\rm d}\sim N(0,\sigma_{\Phi}^{2})$ [32]. However, the true phase-coupled bifrequencies can become indistinguishable from noise when the signal’s power or the strength of the phase coupling is low. This can lead to a false dismissal of PC. Therefore, a test is required to decide whether or not there is phase coupling present at any bifrequency. We assume two hypotheses that are tested on the simulated data:
@@ 4 Application to gravitational-wave data
Importantly, failing to apply artificial phase randomization in the computation of auto-bicoherence can lead to false detection of PC bifrequencies. Consequently, the bicoherence maps also show different patterns depending on whether PR is used or not. One of the blip glitches that occurred in a LIGO Hanford detector during the O2 observation run with high excess power is demonstrated in Figure 4, where the effect of PR can be clearly discerned. We continue to apply PR in all our bicoherence analysis below.
Figure 4: The bicoherence map of a blip glitch, constructed without phase randomization (PR), is shown on the left. The colorbar shows the bicoherence. It incorrectly shows many coupled bifrequencies, which disappear when the bicoherence is computed using phase randomization, as shown in the middle figure. This figure also reveals that the glitch may have been triggered from disturbances in frequencies that up-converted due to phase couplings. Column-3: For comparison, we show here the PR bicoherence map of the background noise data that does not contain any glitch. Such data show the absence of any significant phase coupling. Without the application of PR to this glitch-less data snippet, the bicoherence map (not shown) has numerous points with bicoherence $\approx 1$ strewn across the map.
@@ 4.1 Using bicoherence to classify glitches
Instances of transient excess noise often show up in auxiliary channels and are detected by tools such as Omicron [38] [^1] Also see Ref. [20] for a machine learning based approach to noise transient detection. This tool performs a multi-resolution time-frequency analysis of detector data around such instances. Glitches with excess power above a given threshold are flagged for further inspection in data quality studies and astrophysical searches. Also, the time-frequency (TF) maps or spectrograms of Omicron serve to classify glitches via human intervention and machine learning  [36, 39].
@@
Figure 5: A comparison of second-order (top row) and third-order (bottom row) statistics of glitches. The top row shows the spectrograms of pairs of blips (columns 1-2), koi fish (columns 3-4) and “extremely loud” glitches (columns 5-6). The spectrograms look similar within each category, e.g., blips, simply because the glitches were categorized by the patterns revealed in those maps in the first place. They are different – to varying degrees – across categories, e.g., blips and extremely loud glitches. The bottom row shows the corresponding phase-randomized bicoherence maps, constructed using 4 sec data chunks centered at the glitches. Variations in the bicoherence maps within some glitch categories (e.g., columns 3 and 4 of koi fish) suggest the presence of different phase-coupled bifrequencies or a difference in the nature of that coupling. On the other hand, similarities in these maps across glitch categories (e.g., columns 1 and 4) suggest the presence of similar nonlinear phase couplings in different glitch types. Information in both types of maps, combined, can help in troubleshooting the origin of these noise artifacts.
@@
The bicoherence maps of the blip, koi fish, and extremely loud glitches are produced by applying the phase-randomized auto-bicoherence analysis to look for $3^{\rm rd}$-order similarities in each class. If the observed glitches have a common source, their bicoherence maps will have similarities owing to common bifrequencies related to that source. Such expectations have been borne out in the study of $2^{\rm nd}$-order statistics, such as the TF maps of Omicron, as depicted in the upper panel of Fig. 5. The top row in that figure shows the spectrograms of pairs of blips (columns 1-2), koi fish (columns 3-4) and “extremely loud” glitches (columns 5-6). The spectrograms look similar within each category, e.g., blips, simply because the glitches were categorized by the patterns revealed in those maps. They are different – to varying degrees – across categories (e.g., blips and extremely loud glitches). The bottom row shows the corresponding phase-randomized bicoherence maps, constructed using 4sec data chunks centered at the glitches. Variations in the bicoherence maps within some glitch categories (e.g., columns 3 and 4 of koi fish) suggest the presence of different phase-coupled bifrequencies or a difference in the nature of that coupling. On the other hand, similarities in these maps across glitch categories (e.g., columns 1 and 4) suggest the presence of similar nonlinear phase couplings in different glitch types. (A few more examples of glitches with similar morphology in the third-order statistic are presented in B.) These maps indicate, e.g., that all blips are not made identically. In future, cross-bicoherence maps can help diagnose their origins by revealing what detector systems couple non-linearly to produce them.
It is also interesting to inquire what bicoherence maps look like for data with short-duration Compact Binary Coalescence (CBC) signals. We leave this question for a future work to investigate but provide some preliminary findings in C.
@@ 4.2 Recurrent bifrequencies in the quieter Omicron tiggers
We also note that the nonlinear couplings between the mechanical resonances of test masses reported in our earlier work [23] are found to exist in more recent observation runs as well. Specifically, the bounce and the violin modes of test masses (which register in the bicoherence map at bifrequency (10, 495.2) Hz) as well as the power line harmonics and their sub-harmonics identified in that analysis can be found in O3 too.
@@
198.4, 303.2 501.6 501.61 Violin mode 1st harmonic ITMY Mode 1
501.68 Violin mode 1st harmonic ITMY Mode 2
@@
10.0, 383.2 393.2 393.2 Calibration line non-linearity
16.4, 477.2 493.6 493.61514 Likely calibration line mixing
@@
30.0, 54.4 84.4 84.4 Calibration line non-linearity
38.0, 66.4 104.4 104.4
40.0, 448.8 488.8 488.78625 Likely calibration line mixing
48.8, 250.8 299.6 299.6 Beam splitter violin mode region
@@
68.4, 420.4 488.8 488.78625 Likely calibration line mixing
78.0, 423.6 501.6 501.63223 Violin mode 1st harmonic
@@
134.0, 337.6 471.6 471.68625 Likely calibration line mixing
150.8, 242.4 393.2 393.2 Calibration line non-linearity
155.6, 333.2 488.6 488.78625 Likely calibration line mixing
206.0, 282.4 488.4 488.48624
221.2, 287.6 508.8 508.84515 Violin mode 1st harmonic
Table 2: A subset of gated lines in H1, from O2 and O3, that are associated with high bicoherence. Only some of the bifrequencies with $b_{\rm pr}\geq 0.98$ are listed in column 1, for illustration. The sum of the frequencies in a bifrequency is listed in column 2. Column 3 lists the subset of gated line frequencies in H1, from the same two observation runs, that turn out to be closest to the bifrequency sum listed in column 2. No such association was found in L1 (in O2 and O3) at the same level of significance. It is worth noting that there are additional bifrequencies in those H1 and L1 datasets with equally or more significant bicoherence values whose bifrequency sums do not match any gated line frequencies in that data.